OpenStudy (anonymous):

PLEASE HELP I WILL FAN AND MEDAL!!!!!!!!!!!!!!!! Marneshia walked three over eight of a mile in three over five of an hour. What equation can be used to calculate her unit rate in miles per hour? three eighths over three fifths = five eighths mile per hour three eighths over three fifths = one and three fifths miles per hour three fifths over three eights = one and three fifths miles per hour three fifths over three eights = five eighths mile per hour

OpenStudy (lochana):

did you try it?

OpenStudy (lochana):

distance walked = 3/8 miles time taken = 3/5 hours rate = distance walked/time taken \[= \frac{ \frac{ 3 }{ 8 } }{ \frac{ 3 }{ 5 } }\] 3*5/8*3 = 5/8 three fifths over three eights = five eighths mile per hour

OpenStudy (anonymous):

THANK YOU!!!!!!!!!!!!!!!

OpenStudy (anonymous):

can you help with another?

OpenStudy (lochana):

sure

OpenStudy (anonymous):

Create your own unit rate example in which you have to determine the unit rate of two fractions with different units. Explain what the two fractions are, write the scenario, and calculate the unit rate. Do not forget your units and remember to show your work!

OpenStudy (anonymous):

I'm having sooooooo much trouble with this one

OpenStudy (lochana):

So you can find another situation like your first question to find rate. you can create your own one

OpenStudy (anonymous):

I know but I have no idea what to put

OpenStudy (lochana):

mmm... for example, you can find the rate of water flow into a tank. so it is like, a tank of three over five in size takes nine over four hours to fill completely.

OpenStudy (anonymous):

THANK YOU SOOOOOOOOOOOOOOOOO MUCH!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (lochana):

rate = water volume / time taken rate = (3/5)/(9/4) = 3*4/9*5 rate = 4/15 liters per hour

OpenStudy (lochana):

like that you can come up with anything you like, I think you should create your one.

OpenStudy (anonymous):

alright but thank you so much it really helped!